P is a point on the plane lx + my + nz = p and a point Q is taken on the line such that OP. OQ= p^2 ; show that the locus of point Q is p(lx + my + nz)= x^2 + y^2 + z^2.
1
Expert's answer
2019-09-02T09:19:23-0400

Let P point be "(\alpha,\beta,\gamma)" and Q point be"(x_1,y_1,z_1)"

direction ratios of OP are "\alpha," "\beta," "\gamma" and direction ratios of OQ are "x_1,y_1,z_1."

Since O,Q,P are collinear, we have

"\frac{\alpha}{x_1}=\frac{\beta}{y_1}=\frac{\gamma}{z_1}=k" "......(1)"


As P"(\alpha,\beta,\gamma)" lies on the plane "lx+my+nz=p,"

So,

"l\alpha+m\beta+n\gamma=p"


or "k(x_1+my_1+nz_1)=p" "........(2)"


Given, OP"\times" OQ "=p^2"

"\therefore" "\sqrt{\alpha^2+\beta^2+\gamma^2}\sqrt{(x_1)^2+(y_1)^2+(z_1^2)}=p^2"


or, "\sqrt{k^2(x_1^2+y_1^2+z_1^2)}\sqrt{(x_1)^2+(y_1)^2+(z_1^2)}=p^2"


or, "k(x_1^2+y_1^2+z_1^2)=p^2" "......(3)"


On dividing "(2)" by"(3)" ,we get


"\frac{lx_1+my_1+nz_1}{x_1^2+y_1^2+z_1^2}=\frac{1}{p}"


or, "p(lx_1+my_1+nz_1)=x_1^2+y_1^2+z_1^2"


Hence the locus of point Q is "p(lx+my+nz)=x^2+y^2+z^2."